OpenStudy (anonymous):

Solve: 3x + 7 = 15 - 2x

OpenStudy (anonymous):

add 2x to both sides and subtract 7 from both sides, what does that lead to?

OpenStudy (anonymous):

5x = 4x? :S

OpenStudy (anonymous):

I mea 5x = 8 - 4x

OpenStudy (anonymous):

5x = 8

OpenStudy (anonymous):

The 2x on the right side go out when you put 2x on that side.

OpenStudy (anonymous):

\[3x + 7 = 15 - 2x\]\[3x+2x=15-7\]\[5x=8\]\[x=8/5\]